The following standards for variable manufacturing overhead have been established for a company that makes only one product: Standard hours per unit of output 7.2 hours Standard variable overhead rate $ 13.60 per hour The following data pertain to operations for the last month: Actual hours 2,750 hours Actual total variable manufacturing overhead cost $ 38,090 Actual output 250 units What is the variable overhead efficiency variance for the month?

Multiple Choice

$690 U

$24,480 F

$13,610 U

$12,920 U
Answer #1

D. $12,920 U

Standard hours = Standard hours per unit × Actual output = 7.2 × 250 = 1,800

Variable overhead efficiency variance = Standard rate × (Actual hours − Standard hours) = $13.60 × (2,750 − 1,800) = $12,920 unfavorable